OpenStudy (anonymous):

Choose the ordered pair (x,y) that is a solution of the equation. -8x + 5y = 96

OpenStudy (anonymous):

solve for x first,equate y=0 and solve

OpenStudy (anonymous):

how

OpenStudy (anonymous):

let x=0 first and then solve.

OpenStudy (anonymous):

i need help

OpenStudy (anonymous):

do you have a list of ordered pairs, or are you only supposed to find one?

OpenStudy (anonymous):

A. (-3,11) B. (1,-19) C. (3,19) D. (-2,16)

OpenStudy (anonymous):

i thought so since you have a list, you have to replace \(x\) and \(y\) by the numbers in each ordered pair \((x,y)\) and see if the equality is true or not

OpenStudy (anonymous):

what do you mean

OpenStudy (anonymous):

pick one of the points and i will show you choose one

OpenStudy (anonymous):

c

OpenStudy (anonymous):

(3,19) equation is \(-8x + 5y = 96 \) you replace \(x\) by \(3\) and\(y\) by \(19\) and you get \[-8\times 3+5\times 19=96\] which is either true or false, so we have to compute

OpenStudy (anonymous):

\[-8\times 3+5\times 19=96\] \[-24+95=96\]? no it does not pick another one

OpenStudy (anonymous):

a

OpenStudy (anonymous):

(-3,11) \[-3\times (-3)+5\times 11=96?\] \[9+55=96?\] i don't think so pick another one

OpenStudy (anonymous):

d

jimthompson5910 (jim_thompson5910):

If it was choice d, then... -8x + 5y = 96 -8(-2) + 5(16) = 96 16 + 80 = 96 96 = 96 which is true So (-2,16) does lie on the line, which makes it a solution

OpenStudy (anonymous):

thanks